# A Sparse Tensor Generator with Efficient Feature Extraction

### v2

@@ 4.2 Sparse Tensor Feature Extraction
The features of a sparse tensor have the capacity to reveal the most suitable storage format, partitioning or reordering method that fits best with that tensor if well-examined. The main challenge in extracting the sparse tensor features is to determine which and how many fibers and slices are nonzero. This is because the real-world sparse tensors are highly sparse and contain many empty fibers and slices. One naive approach for sparse tensor feature extraction is to traverse the tensor nonzeros in coordinate (COO) format and update the number of nonzeros of the respective slices and fibers. However, this approach is not practical for large tensors since the real-world tensors often have huge numbers of fibers and slices. For instance, some real tensors have more than $10^{19}$ fibers, and even storing a boolean array of such a large size requires a space of 10 million Terabytes, which is impractical in modern machines. To overcome these challenges, we propose and implement four alternative methods for sparse tensor feature extraction.
@@ 4.2.1 Feature Set
$\mathrm{imbal}_{\mathrm{all}}$ Imbalance (max-$\mathrm{avg}_{\mathrm{all}}$)/max
@@
$\mathrm{imbal}_{\mathrm{nz}}$ Imbalance by excluding empty (max-$\mathrm{avg}_{\mathrm{nz}}$)/max
@@ 4.2.3 Extraction Methods
We propose a hybrid method by combining the sorting- and grouping-based methods. The idea is to utilize different methods for extracting features in different modes within the same tensor. This is because the performance of different methods varies depending on the size of the respective mode. Since we observe that sorting- and grouping-based methods are the best-performing ones at the mode level, we use them interchangeably according to the respective mode size. For mode $m$, we apply the grouping-based method if $size(m)\times size(m+1)<\lambda$ holds, and we apply the sorting-based method otherwise; where $\lambda$ is a predetermined threshold. The reasoning behind this choice is to limit the expected worst-case runtime for the respective modes when exploiting the grouping-based method. We set the threshold $\lambda$ as $10^{9}$ empirically.
@@ 4.3.1 Overview
Our generator considers the significant features of tensors. Consequently, the artificial tensors produced by GenTensor closely emulate real tensors with their respective features and characteristics. A notable advantage of this developed generator is its utilization of size-independent features, such as coefficient of variation, imbalance, and density. This allows for easy generation of instances with varying sizes, enhancing its versatility and applicability in diverse contexts.
The main idea of the proposed tensor generator is to first determine fiber counts per slice, and the number of nonzeros per fiber according to the given coefficient of variation values. The generator has the flexibility to employ any distribution to determine these counts, yet we are currently utilizing normal or log-normal distributions to determine the nonzero layout. The positions of the nonzeros and nonzero slice or fiber indices are selected uniformly. The nonzero values are selected uniformly from the interval (0,1).
@@ 5.1 Experimental setup
The dataset is taken from two real-world sparse tensor collections, namely FROSTT (17) and HaTen2 (16). We have excluded the tensors whose nonzero count is more than 1 billion. We have also excluded the delicious-4d tensor since its fiber count exceeds the maximum value ($1.8\times 10^{19}$) for an unsigned long long int in the C language. As a result, the dataset consists of 16 real-world sparse tensors whose properties are given in Table 3.
@@ 5.2 Performance of Feature Extraction Methods
The performance of four different feature extraction methods in FeaTensor is compared. For this experiment, we use the maximum number of available hardware threads in the machine, which is 96. Since our dataset includes both 3rd-order and higher-order tensors, for a fair comparison, we run all four methods in FeaTensor with the only-3-mode option, i.e. we extract the features corresponding to modes with the largest three sizes. The time measurement covers extracting all features along all modes including the preprocessing time for each method, e.g. sorting time for the sort-based method and time for preparing the hash table for the hash-based method. The experiments validate that all feature extraction methods in FeaTensor give the same and exact results as feature sets, therefore we do not present an accuracy comparison across different methods.
Figure 3 shows the total feature extraction time for each method and each tensor. The tensors are shown in increasing order of nonzero counts. Excluding the hybrid method, we observe that the grouping-based method is superior for smaller tensors, whereas the sort-based method is better for larger tensors. Overall, the hybrid method is the best-performing one for all cases except some small tensors, for which the grouping-based method is the best.
@@
A more detailed mode-level comparison of the sorting- and grouping-based methods is given in Table 4. It is seen that the superiority of a method varies between different modes within the execution of each tensor. Although the grouping-based method tends to be costly in total time due to some modes with larger sizes, it is still the winner for some modes of even large tensors. This is the main reason why integrating the grouping-based method for specific modes within the hybrid approach yields improved performance.
Table 4: Comparison of sort- and group-based methods in terms of mode-level runtime in seconds. Best results for each mode are shown in bold.
Sort-based Group-based
Tensor mode-1 mode-2 mode-3 mode-1 mode-2 mode-3
lbnl-network 0.6 0.4 0.3 0.3 0.6 0.6
nips 1.0 0.7 0.7 0.1 0.1 0.2
uber 1.1 0.7 0.7 0.1 0.1 0.1
chicago-crime-comm 1.8 1.2 1.2 0.1 0.1 0.1
chicago-crime-geo 2.1 1.3 1.3 0.2 0.2 0.2
vast-2015-mc1-3d 8.3 5.7 5.2 0.6 0.8 0.9
vast-2015-mc1-5d 8.0 5.4 5.4 0.8 0.8 1.0
DARPA1998 8.9 6.2 6.7 1.3 103.1 138.0
enron 16.7 14.6 12.4 0.7 0.8 1.2
NELL-2 26.0 20.8 19.1 0.8 0.9 1.5
freebase_music 34.4 28.6 22.0 3.4 16.7 36,000.0
flickr-3d 38.2 29.8 24.6 318.7 4,846.7 13,158.4
flickr-4d 38.9 39.1 24.4 326.7 4,798.0 13,109.2
freebase_sampled 50.6 42.6 32.0 8.3 28.4 36,000.0
delicious-3d 51.1 39.6 31.4 1,209.3 4,840.3 13,517.2
NELL-1 47.4 38.3 33.0 4,346.7 21,478.6 25,526.0
@@ 5.3.1 Feature Quality
In methods that generate values to obey a given density and variation, there is a trade-off between strictly obeying the density or the variation. In other words, if one prioritizes achieving the exact coefficient of variation, the density might get far from the desired value. However, since the number of nonzeros is the most significant feature that a generator must obey for performance concerns, we opted to prioritize adhering to density. For this, in GenTensor we apply some scalings during calculations to catch the given density. This is the reason why the resulting ratios in coefficient of variation seem relatively low compared to the ratios in densities.
@@ 5.3.2 CPD Performance
Although these results demonstrate that GenTensor performs well in most cases, there are a few instances where the generated tensors differ from real tensors in performance. This discrepancy can be attributed to differences in tensor values. Since CPD performance is influenced not only by the sparsity pattern but also by the rank of the tensors—an aspect beyond the scope of this work—these value-related differences explain the observed variations.
@@ 5.3.3 Sensitivity Analysis
To evaluate the sensitivity of our generator, we produce 3 different versions for each generated tensor by taking the seed for pseudo-randomness as 0, 1, and 2. Table 6 shows the nonzero counts and the CPD (SPLATT) runtime for 1 and 48 threads (in seconds). We present the coefficient of variation (CV) for each case. As can be seen in the table, the CV for nonzero count is at most $5\times 10^{-3}$, whereas the CV for CPD performance is at most 0.1. That is, the generator is robust in terms of maintaining a consistent sparsity pattern, and the performance of the generated tensor is not highly sensitive to changes in the seed value. It implies that the proposed generator produces reliable and predictable results, providing a dependable tool for generating sparse tensors for various applications.

### v3

@@ 4.2 Sparse Tensor Feature Extraction
The features of a sparse tensor have the capacity to reveal the most suitable storage format, partitioning or reordering method that fits best with that tensor if well-examined. The main challenge in extracting the sparse tensor features is to determine which and how many fibers and slices are nonzero. This is because the real-world sparse tensors are highly sparse and contain many empty fibers and slices. One naive approach for sparse tensor feature extraction is to traverse the tensor nonzeros in coordinate (COO) format and update the number of nonzeros of the respective slices and fibers. However, this approach is not practical for large tensors since the real-world tensors often have huge numbers of fibers and slices. For instance, some real tensors have more than $10^{19}$ fibers, and even storing a boolean array of such a large size requires a space of 10 million Terabytes, which is impractical in modern machines (See the properties of some real tensors in Table 3). To overcome these challenges, we propose and implement four alternative methods for sparse tensor feature extraction.
@@ 4.2.1 Feature Set
$\mathrm{imbal}_{\mathrm{all}}$ Imbalance $(\text{max}-\mathrm{avg}_{\mathrm{all}})/\text{max}$
@@
$\mathrm{imbal}_{\mathrm{nz}}$ Imbalance by excluding empty $(\text{max}-\mathrm{avg}_{\mathrm{nz}})/\text{max}$
@@ 4.2.3 Extraction Methods
We propose a hybrid method by combining the sorting- and grouping-based methods. The idea is to utilize different methods for extracting features in different modes within the same tensor. This is because the performance of different methods varies depending on the size of the respective mode. Since we observe that sorting- and grouping-based methods are the best-performing ones at the mode level, we use them interchangeably according to the respective mode size. For mode $m$, we apply the grouping-based method if $size(m)\times size(m+1)<\lambda$ holds, and we apply the sorting-based method otherwise; where $\lambda$ is a predetermined threshold. The reasoning behind this choice is to limit the expected worst-case runtime for the respective modes when employing the grouping-based method. We set the threshold $\lambda$ as $10^{11}$ based on empirical evaluations, whose details will be discussed in Section 5.2.
@@ 4.3.1 Overview
Our generator considers the significant features of tensors. Consequently, the arpngicial tensors produced by GenTensor closely emulate real tensors with their respective features and characteristics. A notable advantage of this developed generator is its utilization of size-independent features, such as coefficient of variation, imbalance, and density. This allows for easy generation of instances with varying sizes, enhancing its versatility and applicability in diverse contexts.
The main idea of the proposed tensor generator is to first determine fiber counts per slice, and the number of nonzeros per fiber according to the given coefficient of variation values. The generator has the flexibility to employ any distribution to determine these counts, yet we are currently utilizing normal or log-normal distributions to determine the nonzero layout. The nonzero numerical values are drawn from a uniform (0,1) distribution, while The positions of the nonzeros and nonzero slice or fiber indices are selected uniformly from the corresponding index ranges.
@@ 5.1 Experimental setup
The dataset is taken from two real-world sparse tensor collections, namely FROSTT (17) and HaTen2 (16). We have excluded the tensors whose nonzero count is more than 1 billion. We have also excluded the delicious-4d tensor since its fiber count exceeds the maximum value ($1.8\times 10^{19}$) for an unsigned long long int in the C language. As a result, the dataset consists of 16 real-world sparse tensors whose properties are given in Table 3. Throughout the section, the runtime results are presented as an average of 3 runs.
@@ 5.2 Performance of Feature Extraction Methods
The performance of four different feature extraction methods in FeaTensor is compared. For this experiment, we use the maximum number of available hardware threads in the machine, which is 96. Since our dataset includes both 3rd-order and higher-order tensors, for a fair comparison, we run all four methods in FeaTensor with the only-3-mode option, i.e. we extract the features corresponding to modes with the largest three sizes. The time measurement covers extracting all features along all modes, including the preprocessing time for each method, e.g. sorting time for the sort-based method and time for preparing the hash table for the hash-based method. The experiments validate that all feature extraction methods in FeaTensor give the same and exact results as feature sets, therefore we do not present an accuracy comparison across different methods.
Table 3 presents a detailed runtime comparison between the sort-based and group-based methods at the mode level. The best runtime value for each mode is shown in bold. We also report the value of the decision metric ($size(m)\times size(m+1)$) in the hybrid approach for each mode $m$. The tensors are shown in increasing order of nonzero counts. It is seen that the superiority of a method varies between different modes within the execution of each tensor. Although the grouping-based method tends to be costly in total time due to some modes with larger sizes, it is still the winner for some modes of even large tensors. This is the main reason why integrating the grouping-based method for specific modes within the hybrid approach yields improved performance. We observe that the grouping method yields the best performance when the corresponding decision metric for that mode is less than $10^{11}$ for all cases except LBNL-network, which is the tensor with the smallest nonzero count in our dataset. Therefore, we empirically set the threshold value as $\lambda=10^{11}$ in the hybrid method. This strategy prevents excessive cost in high-dimensional modes and enables faster execution by combining the strengths of both methods.
Table 4: Comparison of sort- and group-based methods in terms of mode-level runtime in seconds, where best results for each mode are shown in bold; along with decision metric ($size(m)\times size(m+1)$) values of hybrid method for each mode $m$, where values less than $\lambda=10^{11}$ are shown in bold.
Sort-based Group-based Decision metric in hybrid method
Tensor mode-1 mode-2 mode-3 mode-1 mode-2 mode-3 mode-1 mode-2 mode-3
LBNL-network 0.60 0.37 0.35 0.32 0.60 0.61 1.8E+07 3.6E+09 3.7E+09
NIPS 1.03 0.69 0.65 0.11 0.14 0.15 7.1E+06 3.5E+07 4.0E+07
uber 1.07 0.68 0.67 0.09 0.10 0.12 2.1E+05 3.1E+05 2.0E+06
chicago-crime-comm 1.75 1.20 1.16 0.11 0.13 0.14 2.5E+03 2.0E+05 4.8E+05
chicago-crime-geo 2.09 1.34 1.30 0.17 0.18 0.19 1.5E+05 2.4E+06 2.4E+06
vast-2015-mc1-3d 8.26 5.71 5.18 0.60 0.82 0.85 2.3E+04 3.3E+05 1.9E+09
vast-2015-mc1-5d 8.00 5.39 5.39 0.78 0.85 1.01 1.1E+06 1.7E+07 1.9E+09
DARPA1998 8.91 6.16 6.73 1.30 103.10 137.98 5.1E+08 5.3E+11 5.3E+11
enron 16.70 14.58 12.37 0.68 0.79 1.19 3.5E+07 1.4E+09 1.5E+09
NELL-2 26.04 20.83 19.07 0.84 0.85 1.51 1.1E+08 2.6E+08 3.5E+08
freebase_music 34.42 28.64 21.99 3.44 16.72 36,000.00 3.9E+09 3.9E+09 5.4E+14
flickr-3d 38.23 29.81 24.63 318.65 4,846.71 13,158.43 5.1E+11 9.0E+12 4.5E+13
flickr-4d 38.87 39.14 24.39 326.67 4,798.01 13,109.18 5.1E+11 9.0E+12 4.5E+13
freebase_sampled 50.63 42.63 32.03 8.31 28.43 36,000.00 2.1E+10 2.1E+10 1.5E+15
delicious-3d 51.05 39.65 31.43 1,209.29 4,840.33 13,517.24 1.3E+12 9.2E+12 4.3E+13
NELL-1 47.36 38.25 33.05 4,346.74 21,478.60 25,525.99 6.2E+12 5.5E+13 7.4E+13
Figure 3 shows the total feature extraction time for each method and each tensor. Excluding the hybrid method, the grouping-based method is superior for smaller tensors, whereas the sort-based method is better for larger tensors. Overall, the hybrid method is the best-performing one for most of the tensors, and it still ranks as the second-best method for the remaining tensors. The results confirm that the hybrid method balances sorting and grouping methods by selecting the appropriate strategy at a mode level according to mode-wise characteristics.
@@ 5.3.1 Feature Quality
In methods that generate values to obey a given density and variation, there is a trade-off between strictly obeying the density or the variation. In other words, if one prioritizes achieving the exact coefficient of variation, the density might get far from the desired value. However, since the number of nonzeros is the most significant feature that a generator must obey for performance concerns, we opted to prioritize adhering to density. For this, in GenTensor we apply some scalings during calculations to catch the given density. It is the reason why the resulting ratios in coefficient of variation seem relatively low compared to the ratios in densities.
@@ 5.3.2 CPD Performance
Although these results demonstrate that GenTensor performs well in most cases, there are a few instances where the generated tensors differ from real tensors in performance. This discrepancy can be attributed to differences in tensor values. Since CPD performance is influenced not only by the sparsity pattern but also by the rank of the tensors, which is an aspect beyond the scope of this work, these value-related differences explain the observed variations. For example, in the case of the chicago-crime-geo tensor, the original nonzeros represent count data and follow a highly skewed distribution. On the other hand, GenTensor assigns nonzeros from a uniform (0,1) distribution. Although structurally consistent, this difference in value distribution may result in different numerical scaling and optimization behavior, which explains the deviation in CPD performance observed with SPLATT. This illustrates how even structurally similar tensors may diverge in runtime depending on their value statistics.
@@ 5.3.3 Sensitivity Analysis
To evaluate the sensitivity of our generator, we produce 3 different versions for each generated tensor by taking the seed for pseudo-randomness as 0, 1, and 2. Table 6 shows the nonzero counts and the CPD (SPLATT) runtime for 1 and 48 threads (in seconds). We present the coefficient of variation (CV) for each case. As can be seen in the table, the CV for nonzero count is at most $5\times 10^{-3}$, whereas the CV for CPD performance is at most 0.1. These results demonstrate that the generator is stable across different random seeds and produces consistent sparsity patterns and decomposition performance. This stability makes GenTensor reliable for repeatable experimentation and benchmarking scenarios.
@@
5.3.4 Robustness and Feature Quality in Higher Orders
To validate the generality of GenTensor, we performed a series of experiments on higher-order tensors with the aim of examining how well the generator preserves target structural properties as tensor dimensionality increases. For this purpose, we generated synthetic tensors with orders $M=6$, $7$, and $8$ where all mode sizes are fixed to 1000. Moreover, to demonstrate the robustness of the generator, we add perturbations in the input structural features (CV and densities) to show they do not drastically change the output structure. Table 7 summarizes the results under six settings, where each row corresponds to a different input configuration.
Table 7: Feature matching and structural robustness results, and average generation times for 6th–8th order tensors generated by GenTensor under six input settings. In all settings, tensor sizes are set to 1000 for all modes. Perturbed Settings 1, 2, 3, 4, and 5 respectively increase $cv_{fib}$, $cv_{nz}$, $d_{slc}$, $d_{fib}$, and $d_{nz}$ target values of “Reference Setting” by 10%, with other features held fixed. The values corresponding to perturbed features for each setting are shown in purple color. The Feature Ratio is the ratio of resulting over target feature values. The Robustness Ratio is the ratio of the feature ratio in the perturbed setting over the feature ratio in the reference setting; and is reported only for the perturbed feature within each setting.
Target Input Features Resulting Generated Features Feature Ratio ( Resulting / Target )
Input Scenario $M$ $cv_{fib}$ $cv_{nz}$ $d_{slc}$ $d_{fib}$ $d_{nz}$ $cv_{fib}$ $cv_{nz}$ $d_{slc}$ $d_{fib}$ $d_{nz}$ $cv_{fib}$ $cv_{nz}$ $d_{slc}$ $d_{fib}$ $d_{nz}$ Robust. Ratio Time (sec)
6 1.0 1.0 1.0E-06 1.0E-08 1.0E-10 0.98 0.98 1.0E-06 9.9E-09 9.9E-11 0.98 0.98 1.00 0.99 0.99 - 1.43
7 1.0 1.0 1.0E-09 1.0E-11 1.0E-13 0.98 0.98 1.0E-09 9.9E-12 9.9E-14 0.98 0.98 1.00 0.99 0.99 - 1.43
Reference Setting 8 1.0 1.0 1.0E-12 1.0E-14 1.0E-16 0.98 0.98 1.0E-12 9.9E-15 9.9E-17 0.98 0.98 1.00 0.99 0.99 - 1.46
6 1.1 1.0 1.0E-06 1.0E-08 1.0E-10 1.08 0.98 1.0E-06 9.9E-09 9.9E-11 0.98 0.98 1.00 0.99 0.99 0.999 1.42
7 1.1 1.0 1.0E-09 1.0E-11 1.0E-13 1.08 0.98 1.0E-09 9.9E-12 9.9E-14 0.98 0.98 1.00 0.99 0.99 0.999 1.43
Perturbed Setting 1 ($cv_{fib}\times$1.1) 8 1.1 1.0 1.0E-12 1.0E-14 1.0E-16 1.08 0.98 1.0E-12 9.9E-15 9.9E-17 0.98 0.98 1.00 0.99 0.99 0.999 1.47
6 1.0 1.1 1.0E-06 1.0E-08 1.0E-10 0.98 1.08 1.0E-06 9.9E-09 9.9E-11 0.98 0.98 1.00 0.99 0.99 0.998 1.44
7 1.0 1.1 1.0E-09 1.0E-11 1.0E-13 0.98 1.08 1.0E-09 9.9E-12 9.9E-14 0.98 0.98 1.00 0.99 0.99 0.998 1.46
Perturbed Setting 2 ($cv_{nz}\times$1.1) 8 1.0 1.1 1.0E-12 1.0E-14 1.0E-16 0.98 1.08 1.0E-12 9.9E-15 9.9E-17 0.98 0.98 1.00 0.99 0.99 0.998 1.48
6 1.0 1.0 1.1E-06 1.0E-08 1.0E-10 0.98 0.98 1.1E-06 9.9E-09 9.9E-11 0.98 0.98 1.00 0.99 0.99 1.000 1.43
7 1.0 1.0 1.1E-09 1.0E-11 1.0E-13 0.98 0.98 1.1E-09 9.9E-12 9.9E-14 0.98 0.98 1.00 0.99 0.99 1.000 1.45
Perturbed Setting 3 ($d_{slc}\times$1.1) 8 1.0 1.0 1.1E-12 1.0E-14 1.0E-16 0.98 0.98 1.1E-12 9.9E-15 9.9E-17 0.98 0.98 1.00 0.99 0.99 1.000 1.48
6 1.0 1.0 1.0E-06 1.1E-08 1.0E-10 1.03 0.98 1.0E-06 1.1E-08 9.9E-11 1.03 0.98 1.00 0.99 0.99 1.000 1.48
7 1.0 1.0 1.0E-09 1.1E-11 1.0E-13 1.03 0.98 1.0E-09 1.1E-11 9.9E-14 1.03 0.98 1.00 0.99 0.99 1.000 1.49
Perturbed Setting 4 ($d_{fib}\times$1.1) 8 1.0 1.0 1.0E-12 1.1E-14 1.0E-16 1.03 0.98 1.0E-12 1.1E-14 9.9E-17 1.03 0.98 1.00 0.99 0.99 1.000 1.54
6 1.0 1.0 1.0E-06 1.0E-08 1.1E-10 0.98 1.03 1.0E-06 9.9E-09 1.1E-10 0.98 1.03 1.00 0.99 0.99 0.999 1.50
7 1.0 1.0 1.0E-09 1.0E-11 1.1E-13 0.98 1.03 1.0E-09 9.9E-12 1.1E-13 0.98 1.03 1.00 0.99 0.99 0.999 1.50
Perturbed Setting 5 ($d_{nz}\times$1.1) 8 1.0 1.0 1.0E-12 1.0E-14 1.1E-16 0.98 1.03 1.0E-12 9.9E-15 1.1E-16 0.98 1.03 1.00 0.99 0.99 0.999 1.58
In our Reference Setting, the target CV parameters are set to $cv_{fib}=cv_{nz}=1.0$; and the target densities are chosen so that the target setting yields approximately 1 million nonzero slices, 10 million nonzero fibers, and 100 million nonzeros for all cases with $M=6$, $7$, and $8$. For instance for $M=6$, it corresponds to assign $cv_{fib}$= $cv_{nz}$=$1.0$, $d_{slc}$=$10^{-6}$, $d_{fib}$=$10^{-8}$, and $d_{nz}$=$10^{-10}$ as input parameters. In each “Perturbed Setting”, a single input feature of “Reference Setting” is increased by 10%, with other features held fixed: Specifically, Perturbed Settings 1, 2, 3, 4, and 5 respectively increase the target values of $cv_{fib}$, $cv_{nz}$, $d_{slc}$, $d_{fib}$, and $d_{nz}$ by 10%. For instance for $M=6$, Perturbed Setting 1 changes $cv_{fib}$ to 1.1; Perturbed Setting 2 changes $cv_{nz}$ to 1.1; Perturbed Setting 3 updates $d_{slc}$ as $1.1\times 10^{-6}$; Perturbed Setting 4 updates $d_{fib}$ as $1.1\times 10^{-8}$; and Perturbed Setting 5 only updates $d_{nz}$ as $1.1\times 10^{-10}$; while keeping other parameters the same as in the Reference Setting. For each setting, we compare the generated structural features against their target values.
For each setting, we compute the ratio between the resulting value over the target input value; to which we refer as the feature ratio. Results indicate that the generated tensors show close agreement with the specified input features across all orders. For example for the Reference Setting, the resulting CV and density values in the generated tensors deviate by less than 2% and 1% from their targets, respectively. Similarly in perturbed settings, GenTensor maintains highly accurate outputs. For all settings and tensor orders, the feature ratios remain between 0.98 and 1.03. These results support that GenTensor can effectively reproduce desired structural characteristics even in higher-dimensional settings.
To evaluate robustness, we compare the generator’s output in the perturbed setting to its output in the reference setting. We define the robustness ratio as the ratio of the feature ratio in the perturbed setting over the feature ratio in the reference setting. In other words, the robustness ratio measures how proportionally the output feature changes in response to a controlled change in the input feature. A robustness ratio close to 1.0 indicates that the generator reacts proportionally to the input perturbation, which is a desirable robustness property. As shown in Table  7, the robustness ratios range from 0.998 to 1.000 for all perturbed settings and tensor orders. These results demonstrate the structural robustness of GenTensor, since it maintains consistent structural behavior under small deviations in input features.
A robustness ratio close to one indicates that the generator reacts proportionally to the input perturbation, which is a desirable robustness property. As shown in Table X, the robustness ratios across all perturbed features and tensor orders range from 0.998 to 1.000. These results demonstrate that the generator is robust, meaning it maintains consistent structural behavior under small deviations in input features.
The last column of Table 7 reports the tensor generation time as an average of three runs. It is seen that perturbing CV or density does not significantly impact runtime. Furthermore, the runtime shows just a slight and expected increase as $M$ increases, indicating that dimensionality alone does not significantly impact the runtime of GenTensor.